OpenStudy (anonymous):

Has an x-intercept of 6 and a 7-intercept of 3 Write an equation in slope-intercept, point-slope, or standard form for the line with the given information. Explain why you chose the form that you used.

OpenStudy (anonymous):

y = 2x -4

OpenStudy (anonymous):

equation starts out in y = mx + b form

OpenStudy (anonymous):

m is the slope, so 2 would be placed before x

OpenStudy (anonymous):

b is the y intercept, so -4 would replace b

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

Yes, thank you so much. I never really understood it c: This helps.

OpenStudy (anonymous):

no problem!

OpenStudy (anonymous):

Mind helping me with another?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Has an x-intercept of 6 and a 7-intercept of 3

OpenStudy (anonymous):

what do you need to do with it? graph it? or find the slope?

OpenStudy (anonymous):

It's the same as "Write an equation in slope-intercept, point-slope, or standard form for the line with the given information. Explain why you chose the form that you used."

OpenStudy (anonymous):

y = mx + b

OpenStudy (anonymous):

the y intercept of 3 is already given to you, so you can replace b with 3

OpenStudy (anonymous):

so you have y = mx + 3, now we just need to figure out the slope

OpenStudy (anonymous):

So it would be y = 6x + 3?

OpenStudy (anonymous):

nah, 6 is the x-intercept, not the slope

OpenStudy (anonymous):

do you know how to graph yet?

OpenStudy (anonymous):

Yes, but I'm still quite confused. I'm quite behind so I'm trying to figure out everything at once.

OpenStudy (anonymous):

okay i'll try to draw it

OpenStudy (anonymous):

you know that x- intercept is 6 and the y intercept is 3

OpenStudy (anonymous):